i have a recipe that calls for 4 ounces of flour, 4 ounces of sugar, 4 ounces of butter, and 3 eggs. using this recipe i should theoretically be able to make 12 cupcakes. if i have 10 ounces of flour, 12 ounces of butter, 14 ounces of Sugar, and 3 eggs how many cupcakes can ACTUALLY make?


1
Expert's answer
2021-02-28T06:19:22-0500

4 ounces of flour + 4 ounces of sugar + 4 ounces of butter + 3 eggs = 12 cupcakes


1 cupcake requires:

4/12=0.33 ounces of flour

4/12=0.33 ounces of sugar

4/12=0.33 ounces of butter

3/12=0.25 eggs


Given amount of ingredients is sufficient for:

10/0.33 = 30.3 cupcakes

12/0.33 = 36.4 cupcakes

14/0.33 = 42.4 cupcakes

3/0.25 = 12 cupcakes


The number of eggs limits the resulting amount of cupcakes. Therefore, a given amount of ingredients is sufficient for 12 cupcakes.